### v2

# Self-Routing RAG: Binding Selective Retrieval with Knowledge Verbalization

Abstract
Selective retrieval improves the accuracy and efficiency of retrieval-augmented generation (RAG) by reducing distractions from low-quality retrievals. However, existing approaches underutilize the inherent knowledge of large language models (LLMs), leading to suboptimal retrieval decisions and degraded generation performance. To bridge this gap, we propose Self-Routing RAG (SR-RAG), a novel framework that binds selective retrieval with knowledge verbalization. SR-RAG enables an LLM to dynamically decide whether to retrieve external knowledge or verbalize its own parametric knowledge. To this end, we design a multi-task objective that jointly optimizes an LLM for knowledge source selection, knowledge verbalization, and response generation. SR-RAG further incorporates a nearest neighbor search mechanism at inference time to improve the accuracy of knowledge source decisions under domain shifts. Fine-tuning three LLMs with SR-RAG significantly improves both their response accuracy and reduces the inference latency. Compared to the strongest selective retrieval baseline, SR-RAG reduces the number of retrievals by 29% while improving performance by 5.1%.
Self-Routing RAG: Binding Selective Retrieval with
Knowledge Verbalization
Di Wuthanks: Equal Contribution, Jia-Chen Gu11footnotemark: 1, Kai-Wei Chang, Nanyun Peng University of California, Los Angeles {diwu,kwchang,violetpeng}@cs.ucla.edu, gujc@ucla.edu
1 Introduction
Retrieval-augmented generation (RAG) equips large language models (LLMs) with external knowledge sources at inference time, enabling stronger performance on tasks requiring up-to-date or domain-specific information (Khandelwal et al., 2020; Lewis et al., 2020; Borgeaud et al., 2022; Ram et al., 2023; Shi et al., 2024). Recently, selective retrieval—an inference strategy that avoids unnecessary retrieval augmentations—has shown promising results in reducing distractions from low-quality retrievals and improving the efficiency of RAG (He et al., 2021; Mallen et al., 2023; Xu et al., 2024; Wu et al., 2024a).
However, a core question has been overlooked by current selective retrieval research:
Does selective retrieval fully honor the knowledge embedded in the LLM itself?
When retrieval is skipped, existing methods use a standard yet simplistic fallback: letting the LLM directly generate the response (Mallen et al., 2023; Jeong et al., 2024; Asai et al., 2024; Wu et al., 2024a). This design prevents the LLM from explicitly articulating its parametric knowledge before generating a response. We argue that the ability to verbalize knowledge, while subtle, has a critical impact on the success of selective retrieval. First, verbalizing internal knowledge expands the LLM’s capacity to answer without retrieval. Prior work demonstrates the ability of LLMs to directly generate high-quality knowledge (Yu et al., 2023) as well as intermediate reasoning paths (Wei et al., 2022; Allen-Zhu and Li, 2025) to benefit the system’s performance. This is especially valuable for complex queries, where naive retrieval methods return irrelevant results and compute-intensive retrieval can return noisy contexts. Second, knowledge verbalization enables more accurate selective retrieval decisions. Existing works train retrieval policies by comparing RAG with direct LLM responses (Wang et al., 2023; Wu et al., 2024a) or resorting to likelihood-based preferences (He et al., 2021; Xu et al., 2024). By contrast, through explicit knowledge elicitation, knowledge verbalization helps characterize the LLM’s capabilities more precisely. Therefore, we argue that knowledge verbalization is a core component of selective retrieval to embrace.
Figure 1: An overview of SR-RAG. Given a user query, the system first selects the most appropriate knowledge source by combining special token prediction with nearest neighbor search. Then, the knowledge is either retrieved from an external source or self-verbalized by the LLM. Finally, the LLM forms the response based on the query and the knowledge. All the steps are streamlined into a single left-to-right generation pass.
We propose Self-Routing RAG (SR-RAG), a selective retrieval framework that tightly integrates knowledge verbalization. By reformulating selective retrieval as a knowledge source selection problem, SR-RAG enables an LLM to self-route between retrieving external knowledge and verbalizing its own parametric knowledge, as illustrated in Figure˜1. Recognizing the limitations of existing frameworks that fine-tune LLMs for selective retrieval via special token prediction (Asai et al., 2024; Wu et al., 2024a), SR-RAG introduces three key innovations. First, we generate diverse verbalizations of the LLM’s internal knowledge to create more accurate training labels for knowledge source selection (Section˜3.3). Second, we introduce a multi-task alignment objective that couples source selection, verbalization, and response generation. We further leverage self-supervised preference alignment over variants of verbalized knowledge to promote high-quality knowledge generation (Section˜3.3). Finally, existing approaches suffer from poor source decision accuracy at inference time due to domain shifts and LLM ability shifts caused by fine-tuning. To bridge this gap, SR-RAG proposes dynamic knowledge source inference via nearest neighbor search, augmenting likelihood-based retrieval decisions with neighboring policy examples in the hidden representation space of the fine-tuned LLM (Section˜3.4). Crucially, SR-RAG’s inference remains efficient, requiring only a single left-to-right generation pass.
We evaluate SR-RAG by fine-tuning Llama-2-7B-Chat (Touvron et al., 2023), Phi-3.5-Mini-Instruct (Abdin et al., 2024), and Qwen2.5-7B-Instruct (Yang et al., 2024) on a mixture of knowledge-intensive tasks to develop the LLMs’ ability to select knowledge sources and improve the quality of knowledge verbalization. Extensive experiments across four benchmarks demonstrate that SR-RAG greatly outperforms both always retrieving and the baseline selective retrieval approach. Compared to the latter, SR-RAG achieves 8.5%/2.1%/4.7% higher overall performance while performing 26%/40%/21% fewer retrievals across the three LLMs respectively (Section˜5.2). Further analyses reveal that SR-RAG improves both the accuracy of selective retrieval decisions (Section˜5.3) and the overall inference efficiency (Section˜5.4). Finally, we conduct thorough ablation studies that confirm the necessity of all three core components to SR-RAG’s strong performance (Section˜5.5). Data and code will be publicly released at https://github.com/xiaowu0162/self-routing-rag.
2 Related Work
Selective Retrieval
To enhance the efficiency of RAG systems and avoid potentially harmful retrievals, several works have proposed selectively skipping retrieval augmentation, which we call selective retrieval following Xu et al. (2024) and Wu et al. (2024a). One popular approach is to assess whether retrieval augmentation increases the likelihood of the LLM generating the correct answer and distill this observation into a supervised decision model (He et al., 2021; Schick et al., 2023; Xu et al., 2024). Analogously, Wang et al. (2023) and Wu et al. (2024a) directly evaluate the correctness of the answers generated with and without retrieval to create the supervision signal. Recent works have also explored solely examining the question’s nature to judge the need for retrieval on question answering tasks (Mallen et al., 2023; Jeong et al., 2024; Asai et al., 2024). To further incorporate the LLM’s confidence in the retrieval decision, Ding et al. (2024), Yao et al. (2024), and Moskvoretskii et al. (2025a) use uncertainty-based selective retrieval. By contrast, this paper highlights the benefits of incorporating knowledge verbalization to make precise selective retrieval decisions and to boost the performance when retrieval is skipped.
Adaptive RAG Inference
This work also relates to the broader field of adaptive RAG, which aims to develop configurable and instance-specific RAG inference strategies. Prior work has introduced active retrieval, where the system dynamically refines or re-issues queries if the initially retrieved content is insufficient (Jiang et al., 2023; Su et al., 2024). Other lines of research explore query decomposition and iterative retrieval to better handle complex questions by breaking them down into manageable sub-queries (Shao et al., 2023; Kim et al., 2023; Liu et al., 2024; Lee et al., 2024). Given the retrieval results, Asai et al. (2024) and Yan et al. (2024) propose inference strategies that critique or revise the retrieved knowledge to improve output quality. Parekh et al. (2025) incorporate an initial decision step to adaptively select the most suitable strategy based on the question. While our work focuses on the more targeted problem of selective retrieval, we introduce a novel form of adaptivity: enabling the LLM to self-route between external retrieval and internal knowledge verbalization, enhancing both flexibility and efficiency in a principled way.
LLMs as Knowledge Sources
A growing body of work has explored using LLMs to generate auxiliary knowledge. Early studies demonstrated the potential of LLMs to produce relevant background knowledge in zero-shot settings for commonsense reasoning (Shwartz et al., 2020; Liu et al., 2022). More broadly, Yu et al. (2023) introduced a generate-then-read approach, treating LLMs as context generators to replace external retrieval for RAG. In addition, LLMs have been shown to generate effective intermediate reasoning steps for complex reasoning questions (Wei et al., 2022; Kojima et al., 2022). Building on these insights, this paper explores fully leveraging LLMs’ knowledge generation ability to benefit selective retrieval systems.
3 Approach
In this section, we first reformulate selective retrieval as a knowledge source selection problem. Then, we introduce the details of the proposed Self-Routing RAG framework.
3.1 Problem Formulation
Knowledge Source-Aware Inference
Given a user query $q$, a knowledge source $s$ is invoked to return relevant information as a text sequence $s(q)$, which is then consumed by an LLM reader $M$ to generate a response $M(q,s(q))$. Standard RAG follows this paradigm, where $s$ typically represents a retriever over an external datastore.
Knowledge Source Selection
In practice, multiple knowledge sources may be available, denoted as $\mathcal{S}=\{\phi,s_{1},...,s_{N}\}$, where $\phi$ is a null knowledge source that returns nothing for any query. A knowledge source selector $P$ chooses the most appropriate source for a given query, i.e., $P(q,\mathcal{S})\in\mathcal{S}$. Selective retrieval can be viewed as the special case where $\mathcal{S}=\{\phi,s\}$ with the full pipeline expressed as $M(q,P(q,\mathcal{S})(q))$.
3.2 Self-Routing RAG: Overview
We propose Self-Routing RAG (SR-RAG), a generalization of selective retrieval that treats the LLM itself as a first-class knowledge source. As illustrated in Figure˜1, given a query $q$, the LLM self-determines which knowledge source to use: either retrieving from an external source or verbalizing its parametric knowledge. The final response is then generated based on both the query and the knowledge collected from the selected source.
Figure 2: Compared to traditional selective RAG, SR-RAG enables an LLM to self-route between knowledge sources and self-act as a knowledge source. We use blue to represent external information and red to represent the LLM and its self-generated tokens.
Building upon traditional selective retrieval methods (Asai et al., 2024; Wu et al., 2024a), SR-RAG fine-tunes the LLM to streamline its inference process with special tokens, enabling efficient inference with a single left-to-right generation pass. Three sets of special tokens are introduced:
  1. 1.
    <EOQ> marks the end of the query and prompts the LLM for knowledge source selection [^1] This design aligns with Wu et al. (2024a) but diverges from Asai et al. (2024). We argue that this token is necessary for the LLM to allocate probability mass to the tokens for the knowledge sources..
  2. 2.
    A set of special tokens <s>, each representing a knowledge source $s$. Our main setup uses two sources: external ($S_{e}$) and internal ($S_{i}$). This formulation can naturally accommodate more than two knowledge sources as well.
  3. 3.
    <EOK> indicates the end of the knowledge and triggers the LLM to generate the answer.
As shown in Figure˜2, compared to traditional selective retrieval, SR-RAG enables the LLM to actively select the best knowledge source and seamlessly act as a knowledge source itself.
3.3 Self-Routing RAG: Training
To train the backbone LLM for SR-RAG, we propose a pipeline to mine self-supervision from widely available question answering or instruction-following data, consisting of pairs of question $q$ and response $a$. Our pipeline uses the LLM itself as the internal knowledge source $S_{i}$, alongside the external knowledge source $S_{e}$, without requiring additional human supervision or synthetic labels from stronger LLMs, demonstrating strong scalability.
Data Construction
To enable an LLM to accurately determine whether a question falls within its parametric knowledge and to robustly elicit the knowledge, we argue that thorough and diverse knowledge verbalization is crucial. Following this intuition, SR-RAG collects contexts by exploring each of the available knowledge sources:
  • Parametric Knowledge Verbalization: We leverage GenRead (Yu et al., 2023) to elicit knowledge from the LLM parametric knowledge source $S_{i}$ and generate $n$ diverse verbalized contexts, denoted as $c_{i_{1}},c_{i_{2}},...,c_{i_{n}}$.
  • External Knowledge Retrieval: We retrieve $n$ context chunks from the external knowledge source $S_{e}$, denoted as $c_{e_{1}},c_{e_{2}},...,c_{e_{n}}$. In this work, we consider retrieving from Wikipedia with an off-the-shelf dense retriever.
Then, each context $c_{j}$ is scored using the log-likelihood $l_{j}=p_{M}(a|q,c_{j})$ of generating the correct answer. Based on the ranking of $l_{j}$, we identify the preferred source $s\in\{S_{i},S_{e}\}$ as the one contributing to the majority of the top-$n$ ranked contexts. The resulting $(q,a,s,\{c_{j},l_{j}\})$ tuples are saved for model training. For convenience of later reference, we denote the contexts from $S_{i}$ and $S_{e}$ that lead to the highest and lowest likelihoods as $c_{i+},c_{i-},c_{e+},c_{e-}$, respectively. We present the formal data creation algorithm in Section˜B.1.
Objective
SR-RAG proposes a two-stage multi-task learning framework that jointly optimizes knowledge source selection, knowledge verbalization, and response generation. The first stage performs behavior cloning on three losses:
  1. 1.
    $\mathcal{L}_{src}$: a cross-entropy loss for the preferred knowledge source $s$ following <EOQ>:
    $\mathcal{L}_{src}=-\log p_{M}(\texttt{\textless s\textgreater}|q),$ (1)
    where <s> represents the actual token corresponding to the chosen source $s\in\mathcal{S}$.
  2. 2.
    $\mathcal{L}_{verb}$: a cross-entropy loss on the knowledge tokens, only when the LLM itself ($S_{i}$) is labeled as the preferred knowledge source:
    $\mathcal{L}_{verb}=\begin{cases}-\log p_{M}(c_{i+}|q),&\text{if }s=S_{i},\\ 0,&\text{if }s=S_{e}.\end{cases}$ (2)
  3. 3.
    $\mathcal{L}_{ans}$: a cross-entropy loss on generating the answer based on the query and the knowledge from the preferred source:
    $\mathcal{L}_{ans}=\begin{cases}-\log p_{M}(a|q,c_{i+}),&\text{if }s=S_{i},\\ -\log p_{M}(a|q,c_{e+}),&\text{if }s=S_{e}.\end{cases}$ (3)
The final loss for the first stage is a simple combination of the three objectives:
$\mathcal{L}_{stage1}=\mathcal{L}_{src}+\mathcal{L}_{verb}+\mathcal{L}_{ans}.$ (4)
To further boost the LLM’s ability to generate useful knowledge, SR-RAG incorporates a second-stage fine-tuning via direct preference optimization (DPO) (Rafailov et al., 2023), pairing self-verbalized knowledge with self-generated preference labels $(c_{i+},c_{i-})$.
$\mathcal{L}_{stage2}=\mathcal{L}_{src}+\mathcal{L}_{verb}^{DPO}+\mathcal{L}_{ans},$ (5)
$\mathcal{L}_{verb}^{DPO}=\begin{cases}-\log\sigma\Big(\beta\log\frac{p_{M}(c_{i+}|q)}{p_{ref}(c_{i+}|q)}-\beta\log\frac{p_{M}(c_{i-}|q)}{p_{ref}(c_{i-}|q)}\Big),&\text{if }s=S_{i},\\ 0,&\text{if }s=S_{e}.\end{cases}$
$M$ and refref are initialized with the LLM fine-tuned on $\mathcal{L}_{stage1}$, and only $M$ is updated.
Overall, this self-supervised pipeline effectively binds knowledge verbalization with the selective retrieval paradigm, enabling the LLM to learn accurate knowledge source preferences through performance-oriented labeling. Analogous to distilling complex “System 2” reasoning into fast “System 1” inference (Yu et al., 2024), the DPO objective leverages the computationally expensive high-quality knowledge to teach the LLM cost-efficient knowledge verbalization at inference time. Finally, SR-RAG naturally extends to more than two knowledge sources (further discussions and experiments in Appendix˜C), which is useful for distinguishing domain-specific corpora or retrieval methods with varying cost-quality trade-offs.
3.4 Self-Routing RAG: Inference
As shown in Figures˜1 and 2, SR-RAG inference unfolds in a single left-to-right pass through three steps: source selection, knowledge collection, and answer generation.
Nearest Neighbor-Enhanced Source Selection
A common approach to selecting a knowledge source is to compare the likelihood $p_{M}(\text{{\textless s\textgreater}}|q)$ for each $s\in\mathcal{S}$, against a fixed threshold (Asai et al., 2024; Wu et al., 2024a). However, this approach fails to account for shifts in the LLM’s ability after fine-tuning and lacks fine-grained control over the decision boundary. To make source selection more robust, we propose a dynamic nearest neighbor-based source selection mechanism that builds a policy datastore based on rollouts after the fine-tuning. Concretely, the fine-tuned LLM is first evaluated on a set of question-answer pairs [^2] Empirically, we reuse the training set so that no additional data or supervision is required.. For each pair, the probabilities of generating the answer $a$ conditioned on different knowledge sources are compared to decide the preferred source. We then build a policy datastore mapping each query to its preferred source, using the hidden representation at <EOQ> as the key. At inference, we retrieve $k$ nearest neighbors from the policy datastore and use their source labels to form a distribution over the sources $p_{D}(\text{{\textless s\textgreater}}|q)$. Finally, to select the best source $s\in\mathcal{S}$, we apply a threshold on the product:
$p_{M}(\text{{\textless s\textgreater}}|q)\times p_{D}(\text{{\textless s\textgreater}}|q).$ (7)
While tackling the challenges of source selection due to shifts in the LLM’s ability, this approach also exhibits better interpretability. Since the policy datastore consists of explicit source assignments, it can be audited, modified, and expanded by human experts to steer SR-RAG’s retrieval behavior in a fine-grained manner.
Subsequently, the knowledge from the corresponding source is gathered. If the LLM prefers $S_{i}$, we use greedy decoding to directly verbalize a single knowledge context. After SR-RAG fine-tuning, the generated context serves as a compressed yet high-quality articulation of the parametric knowledge, which would otherwise require compute-expensive knowledge verbalization to elicit. If an external source <s> is selected instead, we pause the decoding and retrieve from knowledge source $s$. With the verbalized or retrieved knowledge appended to the context, the LLM proceeds to generate the final response.
Figure 3: Knowledge verbalization significantly affects the LLM ability boundary. For a large number of instances (16.4% - 38.8%, orange), GenRead reverses the knowledge source preferences: without considering GenRead, RAG dominates over parametric knowledge.
4 Experimental Setup
4.1 Implementation Details of SR-RAG
Data Construction
The main experiments are performed on two knowledge sources [^3] We present multi-source extensions of SR-RAG in Appendix C.: the 2018 English Wikipedia (<Wiki>) as the external knowledge source, and the LLM itself (<Self>) as the internal knowledge source. We use the official Wikipedia embeddings released by Karpukhin et al. (2020) and retrieve at a granularity of 100-word chunks. GenRead (Yu et al., 2023) is used to verbalize diverse knowledge contexts. GenRead clusters zero-shot knowledge verbalizations in as in-context demonstrations to verbalize diverse knowledge. We limit the verbalized knowledge chunks to a maximum of 150 tokens. From each knowledge source, we collect $n$ = 5 knowledge contexts.
Training
We fine-tune on a mixture of six short- and long-form knowledge-intensive datasets: Wizard of Wikipedia (Dinan et al., 2019), Natural Questions (Kwiatkowski et al., 2019), FEVER (Thorne et al., 2018), OpenBookQA (Mihaylov et al., 2018), ARC-Easy (Bhakthavatsalam et al., 2021), and ASQA (Stelmakh et al., 2022). This mixture of 53,042 instances is a subset of the RAG instruction tuning data shown effective in Asai et al. (2024). After running the data construction algorithm, 46.9% of the instances are labeled with <Self> and the rest are labeled with <Wiki> as the preferred knowledge source, averaged across three LLMs. Full training details are provided in Section˜B.1.
Inference
[…]
4.2 Evaluation
Datasets and Metrics
We evaluate on a diverse set of four knowledge-intensive NLP tasks. PopQA (Mallen et al., 2023) is a free-formed long-tail open-domain question answering dataset. Following Asai et al. (2024), we use the subset of 1,399 questions that aims to test knowledge of long-tail entities. TriviaQA (Joshi et al., 2017) is an established open-domain question answering dataset that features relatively complex and diverse questions. We use the same test split and retrieval setup as in Asai et al. (2024). PubHealth (Zhang et al., 2023) is a fact-checking dataset focusing on checking claims in the public health domain. Finally, ARC Challenge (Bhakthavatsalam et al., 2021) is a multiple-choice question answering dataset featuring grade-school level science questions. Following common practice, we perform lexical post-process of the model’s output and report accuracy for PubHealth and ARC and substring matching for PopQA and TriviaQA.
Baselines
We compare SR-RAG with baselines that cover various training and inference strategies. (1) First, using the LLM before fine-tuning, we compare with either always retrieving or always verbalizing with GenRead. (2) As illustrated in Figure˜2, our main baseline is the state-of-the-art prior selective retrieval pipeline, combining the advantage of He et al. (2021), Asai et al. (2024), and Wu et al. (2024a). Specifically, the likelihoods of the LLM generating the answer with and without retrieval are used to create the knowledge source selection label. Then, we fine-tune the LLM for knowledge source selection (among $S_{e}$ and $\phi$) and generate the answer with optional retrieval. At inference, we apply a uniform threshold of 0.2 to the likelihood of the retrieval token for selective retrieval. (3) Always retrieving with the fine-tuned SR-RAG LLM.
5 Results
5.1 Knowledge Verbalization Alters Knowledge Source Preference
Training Inference PopQA TriviaQA PubHealth ARC Average
ACC %RAG ACC %RAG ACC %RAG ACC %RAG ACC %RAG
Llama-2-7B-Chat
No Fine-tuning Always RAG 0.529 100% 0.641 100% 0.457 100% 0.546 100% 0.543 100%
GenRead 0.247 0% 0.616 0% 0.515 0% 0.605 0% 0.496 0%
Selective RAG Always RAG 0.567 100% 0.640 100% 0.588 100% 0.588 100% 0.596 100%
Selective RAG 0.565 98% 0.638 100% 0.589 100% 0.594 65% 0.597 86%
SR-RAG Always RAG 0.568 100% 0.669 100% 0.689 100% 0.608 100% 0.634 100%
SR-RAG 0.566 96% 0.664 89% 0.730 40% 0.630 29% 0.648 64%
Phi-3.5-Mini-Instruct
No Fine-tuning Always RAG 0.541 100% 0.594 100% 0.549 100% 0.771 100% 0.614 100%
GenRead 0.331 0% 0.567 0% 0.442 0% 0.840 0% 0.545 0%
Selective RAG Always RAG 0.570 100% 0.645 100% 0.701 100% 0.813 100% 0.682 100%
Selective RAG 0.570 100% 0.638 95% 0.704 91% 0.815 83% 0.682 92%
SR-RAG Always RAG 0.567 100% 0.659 100% 0.689 100% 0.820 100% 0.684 100%
SR-RAG 0.566 98% 0.657 92% 0.705 24% 0.854 5% 0.696 55%
Qwen2.5-7B-Instruct
No Fine-tuning Always RAG 0.563 100% 0.667 100% 0.446 100% 0.916 100% 0.648 100%
GenRead 0.334 0% 0.626 0% 0.676 0% 0.875 0% 0.628 0%
Selective RAG Always RAG 0.555 100% 0.654 100% 0.600 100% 0.827 100% 0.659 100%
Selective RAG 0.529 88% 0.648 93% 0.608 82% 0.835 78% 0.655 85%
SR-RAG Always RAG 0.573 100% 0.662 100% 0.596 100% 0.821 100% 0.663 100%
SR-RAG 0.572 99% 0.659 89% 0.682 34% 0.830 46% 0.686 67%
Table 1: Main evaluation results on four tasks. The best results are boldfaced. Across three backbone LLMs, SR-RAG consistently outperforms selective RAG and always retrieving while using a much lower retrieval budget.
To motivate our approach, we first show that knowledge verbalization substantially impacts on identifying when an LLM requires retrieval. Figure˜3 presents a pilot study on Llama-2-7B-Chat using four datasets from SR-RAG’s training mixture. For each instance, we compute the likelihood of the LLM generating the correct answer with no context (blue), with the most helpful GenRead verbalization ($c_{i+}$, orange), and with the most helpful retrieved passage ($c_{e+}$, red). Notably, GenRead changes the preferred knowledge source in 16% of Natural Questions examples and over 30% in the other datasets. This suggests that prior selective retrieval methods, which omit verbalization, may significantly underestimate the LLM’s capabilities, reinforcing the need to incorporate knowledge verbalization for accurate source preference labeling.
Method PopQA TriviaQA PubHealth ARC Average
Accuracy (Verbalization $\geq$ Retrieval)
Self-RAG 0.957 0.936 0.867 0.908 0.917
SR-RAG w/o. kNN 0.959 0.930 0.869 0.888 0.912
SR-RAG 0.959 0.943 0.880 0.910 0.923
AUROC (Retrieval >> Verbalization)
Self-RAG 0.489 0.503 0.438 0.557 0.497
SR-RAG w/o. kNN 0.490 0.567 0.564 0.513 0.534
SR-RAG 0.577 0.565 0.606 0.533 0.570
Table 2: Source selection accuracy measured on Llama-2-7B-Chat. SR-RAG achieves the best averaged performance in both evaluation settings.
Figure 4: Accuracy and system latency of SR-RAG fine-tuned Llama-2-7B-Chat with different verbalization frequencies. SR-RAG’s source selection policy (marked with stars) achieves near-optimal accuracy-efficiency trade-off without dataset-specific thresholds.
Training Inference PopQA TriviaQA PubHealth ARC Average
ACC %RAG ACC %RAG ACC %RAG ACC %RAG ACC %RAG
SR-RAG SR-RAG 0.566 96% 0.664 89% 0.730 40% 0.630 29% 0.648 64%
SR-RAG SR-RAG w/o. kNN 0.558 94% 0.658 77% 0.694 72% 0.627 56% 0.634 75%
SR-RAG w/o. kv. label SR-RAG 0.568 100% 0.644 100% 0.598 100% 0.629 84% 0.610 96%
SR-RAG w/o. $\mathcal{L}_{verb}^{DPO}$ SR-RAG 0.564 98% 0.645 100% 0.674 100% 0.581 66% 0.616 86%
Table 3: Ablation studies on SR-RAG. Llama-2-7B-Chat is used as the LLM.
5.2 SR-RAG Improves Accuracy with Reduced Retrieval Frequency
Table˜1 shows the end-to-end generation performance on three LLMs, demonstrating the advantage of SR-RAG over always retrieving, selective retrieval, and other baselines. Relative to always retrieving with the pretrained model, SR-RAG improves accuracy by 19.3% (Llama-2-7B-Chat), 13.4% (Phi-3.5-Mini-Instruct), and 5.9% (Qwen2.5-7B-Instruct). Against the strongest selective retrieval baseline, SR-RAG yields 8.5%/2.1%/4.7% higher accuracy. Interestingly, despite saving 8–15% retrievals, the selective retrieval baseline achieves nearly identical accuracy to always retrieving. This suggests that selective RAG alone cannot reliably identify queries that do not require retrieval. In contrast, SR-RAG reduces 20–40% retrievals while maintaining or improving performance thanks to accurate knowledge source selection and high-quality verbalization.
Remarkably, with a uniform inference datastore and threshold, SR-RAG dynamically adapts its retrieval behavior based on dataset characteristics. For a dataset that emphasizes long-tail knowledge like PopQA, SR-RAG tends to retrieve external knowledge most of the time. On the other hand, for PubHealth and ARC where the model’s knowledge may suffice for a number of questions, SR-RAG relies on internal knowledge more confidently, resulting in much better performance compared to always retrieving. Qualitative examples illustrating SR-RAG’s behavior can be found in Figure˜7 and Figure˜8 in the appendix.
5.3 SR-RAG Makes Accurate Source Selection Decisions
Can SR-RAG accurately choose the correct knowledge source? Table˜2 compares SR-RAG with Self-RAG (Asai et al., 2024) and an SR-RAG variant without kNN-based policy. We evaluate accuracy using two criteria: (1) abstaining from retrieval when it does not harm performance (top) and (2) retrieving only when it provides strictly better context (bottom). While all methods perform well under the first criterion, SR-RAG achieves the highest overall accuracy. Under the second, stricter definition, SR-RAG demonstrates substantially better performance, outperforming Self-RAG by 14.7% in average AUROC. Removing kNN-based policy (w/o. kNN) leads to large drops in both accuracy and AUROC, highlighting the benefit of adapting to fine-tuning-induced shifts in LLM ability.
5.4 System Efficiency
We assess SR-RAG’s end-to-end latency using Llama-2-7B-Chat under batched inference [^6] Detailed latency modeling formulation in Section B.2. In our setting, latency is linearly correlated with retrieval frequency under small batch sizes.. Figure˜4 shows the trade-off between accuracy and latency under varying retrieval proportions. As expected, latency improves as fewer instances invoke retrieval. Due to dataset-specific difficulty, the optimal retrieval proportion varies per dataset. However, SR-RAG’s learned source selection policy achieves a near-optimal accuracy-efficiency trade-off across all datasets, without any dataset-specific threshold tuning. This confirms that SR-RAG not only improves performance but does so efficiently.
5.5 Ablation Study
To understand the contribution of each component, we perform ablations on Llama-2-7B-Chat (Table˜3). Removing kNN-based source selection (w/o. kNN) results in lower accuracy and higher retrieval rates, confirming that kNN inference helps the model adapt to ability shifts introduced by fine-tuning. Disabling knowledge verbalization during preference labeling (w/o. kv. label) causes the LLM to over-rely on retrieval and reduces performance, showing that verbalization is crucial for effective labeling. Finally, $\mathcal{L}_{stage2}$ is ablated and the $\mathcal{L}_{stage1}$ loss is kept for stage 2 training. The result shows significantly more retrievals as the model’s knowledge verbalization ability degrades. These findings highlight that all components are vital to SR-RAG’s success. In Appendix˜D, we further analyze the <EOQ> representations, other SR-RAG hyperparameters, as well as qualitative examples.
6 Conclusion
We present SR-RAG, a novel retrieval-augmented generation (RAG) framework that tightly integrates selective retrieval with knowledge verbalization. By reformulating selective retrieval as a knowledge source selection problem, SR-RAG enables the LLM to not only choose between external and internal knowledge sources but also to serve as a knowledge source itself. During inference, SR-RAG leverages internal hidden states and a nearest-neighbor policy to make accurate, adaptive source selection decisions. Extensive experiments show that SR-RAG significantly improves answer accuracy while reducing retrieval frequency and latency, offering a scalable and reliable path forward for more efficient, knowledge-aware RAG systems.
Supplementary Material: Appendices
Appendix A List of Notations
In Table˜4, we present the major notations and parameters used throughout the paper.
Notation Description
$q$ User query input to the system.
$a$ The expected answer.
$M$ The LLM.
$\mathcal{S}$ Set of all knowledge sources.
$S_{e}$ The external knowledge source.
$S_{i}$ The internal knowledge source (parametric knowledge).
$c_{i+}$ Most helpful verbalized knowledge context from $S_{i}$.
$c_{i-}$ Least helpful verbalized knowledge context from $S_{i}$.
$c_{e+}$ Most helpful retrieved knowledge context from $S_{e}$.
$c_{e-}$ Least helpful retrieved knowledge context from $S_{e}$.
<EOQ> End-of-query special token.
<EOK> End-of-knowledge special token.
<s> Special token representing knowledge source $s\in\mathcal{S}$.
<Wiki> Special token representing Wikipedia.
<Self> Special token representing the LLM itself as knowledge source.
$k$ number of neighbors retrieved for source policy inference
Table 4: A summary of the key symbols and parameters used in the paper.
Appendix B SR-RAG: Further Details
B.1 Training Details
Dataset Construction Algorithm
Algorithm˜1 presents the full algorithm for constructing training data and labeling knowledge source preferences. GenRead is executed independently on each training data subset. We adopt instance-level notation for clarity. The pipeline naturally scales to additional knowledge sources by applying knowledge collection and likelihood evaluation in parallel across sources.
Algorithm 1 SR-RAG Training Data Construction
1:LLM $M$, External Retriever $\mathcal{R}$, Dataset $\mathcal{D}$, Number of contexts $n$
2:for $(q,a)\in\mathcal{D}$ do
3:   // Retrieving External Knowledge
4:   $\mathcal{C}_{S_{e}}\leftarrow\mathcal{R}(q,n)$
5:   // Knowledge Verbalization
6:   $\mathcal{C}_{S_{i}}\leftarrow$ $M$.GenRead(q,n)(q,n)
7:   // Compute Likelihoods
8:   for $c\in\mathcal{C}_{S_{e}}\cup\mathcal{C}_{S_{i}}$ do
9:     $l_{c}\leftarrow p_{M}(a|q,c)$
10:   end for
11:   $s\leftarrow\arg\max_{s\in\{S_{e},S_{i}\}}\sum_{c\in\mathcal{C}_{s}}l_{c}$
12:   Store $(q,a,s,\{c,l_{c}\})$
13:end for
14:return Processed dataset with labeled knowledge sources
GenRead Prompt
We implement GenRead following the setup in Yu et al. (2023). In the second verbalization round, five clusters with five in-context examples each are used. All datasets except ASQA follow the general prompt shown in Figure˜5. For ASQA, we include an additional instruction to handle ambiguity: “If the question is ambiguous, generate multiple documents for each possibility.
Generate a background document from Wikipedia to help answer the following question. Directly start with document content and do not generate URL. Question: {question} Background document:
Figure 5: Prompt used for knowledge verbalization data collection via GenRead.
Training Data
Table˜5 summarizes the training and validation splits, including the proportion of examples where self-verbalized knowledge is preferred over retrieved knowledge.
Dataset Train Validation Total %Verbalization
Llama Phi Qwen
ARC_Easy 2037 107 2144 61% 84% 66%
NQ 14753 776 15529 28% 33% 41%
OBQA 4462 234 4696 61% 77% 61%
FEVER 9467 498 9965 52% 58% 68%
WoW 16493 868 17361 13% 55% 32%
ASQA 3700 194 3894 13% 25% 16%
Table 5: Statistics of the training and validation data with verbalization percentages. Llama = Llama-2-7B-Chat, Phi = Phi-3.5-Mini-Instruct, and Qwen = Qwen2.5-7B-Instruct.
Training Process
To fully leverage the backbone LLM’s ability to follow natural language instructions, both SR-RAG fine-tuning and inference use the following prompt that interleaves special tokens with natural language:
Question: {question} Background knowledge: <EOQ>
<s> {knowledge} <EOK> Answer: {answer}
During training, the loss is computed on the knowledge part only if <s> is <Self>. If <s> is <Wiki>, we augment $c_{e+}$ by randomly appending $\max(p-1,0)$ retrieved contexts, where $p$ is sampled from a Poisson distribution with $\lambda=2$. This data augmentation improves the LLM’s robustness to different retrieval strategies and various levels of retrieval quality. For stage 1 training, we use batch size 64, learning rate 1e-5, and fine-tune for 1 epoch. For stage 2 training, we use batch size 64, learning rate 5e-7, $\beta$ = 0.3 for DPO, and train for another epoch. All the experiments are performed on a machine with eight A800 (80GB) GPUs and a machine with eight A6000 GPUs. On eight A800 (80GB) GPUs, the two-staged training takes approximately 10 hours for a 7B-sized model.
B.2 Latency Formulation
To evaluate the inference efficiency of SR-RAG, we measure the latency in a realistic batched inference setup, where the system handles a batch of $B$ = 10 queries and returns the results for all queries. For the latency experiments presented in the paper, we decompose the system latency as follows:
  • Source Selection Time ($T_{d}$): The time for the knowledge source selector to determine whether to retrieve from external sources or rely on parametric knowledge.
  • Retrieval Latency ($T_{rs}$): The time for fetching external knowledge from the database if the model chooses to retrieval from an external knowledge source $s$. In our batched setting, we calculate $T_{rs}$ by performing a batched retrieval for all the $B$ instances that require retrieval and report the per-item latency.
  • Verbalization Latency ($T_{v}$): The time for the LLM to verbalize parametric knowledge if itself is selected as the knowledge source.
  • Generation Latency ($T_{g}$): The time for the LLM to generate the response, conditioned on either retrieved or verbalized knowledge.
Thus, the total per-item latency $T_{\text{total}}$ is given by:
$T_{\text{total}}=\begin{cases}T_{d}+T_{v}+T_{g},&\text{if verbalize},\\ T_{d}+T_{rs}+T_{g},&\text{if retrieve from $s$}.\end{cases}$
We have the following remarks:
  • This formulation assumes that both the retrieval index and the source selection datastore are pre-constructed. This assumption is reasonable as these indices are only constructed once and then constantly reused.
  • We choose the batched setup due to the complexity of the retrieval system. For instance, in our implementation of Wikipedia search, it takes around five to ten seconds per instance to encode the query and retrieve the most relevant context chunks. The batched setting amortizes the latency of single retrieval.
  • As the nearest neighbor only involves one matrix product for similarity calculation and one top-k operation, $T_{d}$ is generally very small. In fact, we find $T_{d}$ (0.01s) <<<< $T_{g}$ (0.1s) << $T_{v}$ (1s). On the other hand, $T_{rs}$ is the major bottleneck of the pipeline. As a result, in an online setting, the system’s efficiency gain directly converges to the percentage of retrievals it is able to avoid.
Appendix C SR-RAG: Multi-Source Extensions
Beyond routing between the LLM itself and an external knowledge source, SR-RAG’s formulation is naturally compatible with routing between multiple external knowledge sources as well. In this section, we provide further experiments on routing between three knowledge sources and discuss other potential multi-source applications.
C.1 Experiment: Routing between General and Specialized Sources
We perform experiments on the PubMed corpus (National Center for Biotechnology Information (), NCBI), which contains professional biomedical literature. Compared to Wikipedia and the LLM itself, PubMed has the unique advantage of featuring long-tail specialized biomedical knowledge. Meanwhile, it falls short at providing knowledge for many general factuality questions, posing a challenge to accurately routing the knowledge source.
Based on this new setup, we formulate SR-RAG with three knowledge source tokens: <Self>, <Wiki>, and <Pubmed>. We leverage the MedRAG toolkit (Xiong et al., 2024) to implement retrieval on PubMed. The training and inference strategies of the two-source SR-RAG recipe naturally extends to this formulation and are followed in our experiments here. For training and evaluation, we additionally use PubMedQA (Jin et al., 2019), a widely used biomedical question answering dataset. PubMedQA features short questions extracted from PubMed abstracts and provides associated ground truth yes/no/maybe answers. We sample 10k artifically generated documents for training (combined with the data mixture used in the main text) and 500 examples from the expert labeled subset for evaluation. The same set of LLMs from three families (Llama, Qwen, Phi) are trained and evaluated.
In Table˜6, we present the three-way routing results of SR-RAG and compare with always RAG as well as GenRead with the model before fine-tuning. Our main findings are as follows:
  • Adding PubMed as an additional knowledge source does not harm its effectiveness. The results on datasets other than PubMedQA aligns with the performance of two-source SR-RAG reported in the main text.
  • SR-RAG training improves the model’s ability to verbalize its knowledge on PubMedQA as well while being much more efficient than GenRead, as shown by the better performance of SR-RAG + Always Verbalizing compared to GenRead.
  • SR-RAG intelligently routes the best knowledge source for different workloads. For PubMedQA, PubHealth, and ARC, where PubMed can outperform Wiki by a large margin, SR-RAG frequently self-selects PubMedQA as its preferred knowledge source. On the other hand, on general factuality benchmarks such as PopQA and TriviaQA, SR-RAG selects Wikipedia more frequently. More importantly, SR-RAG achieves this behavior end-to-end without the need for any manual interventions.
  • On PubMedQA, the end-to-end SR-RAG achieves the best performance compared to the baselines. The trend is consistent across three different LLMs.
Overall, this experiments confirms the effectiveness of extending SR-RAG to three sources and suggests a promising outlook for further extension.
Training Inference PubMedQA PopQA TriviaQA PubHealth ARC Average
Llama-2-7B-Chat
No Fine-tuning GenRead 0.544 0.247 0.616 0.515 0.605 0.505
Always RAG, Wiki 0.636 0.529 0.641 0.457 0.546 0.562
Always RAG, PubMed 0.656 0.157 0.470 0.612 0.537 0.486
SR-RAG (3-source) Always RAG, Wiki 0.342 0.587 0.639 0.668 0.639 0.572
Always RAG, PubMed 0.672 0.325 0.539 0.713 0.645 0.642
Always Verbalizing 0.388 0.305 0.572 0.724 0.631 0.579
SR-RAG 0.680 0.583 0.647 0.724 0.655 0.677
(% <Wiki>) 6% 95% 88% 17% 4% 29%
(% <Pubmed>) 91% 4% 7% 70% 86% 64%
(% <Self>) 3% 1% 5% 13% 10% 8%
Phi-3.5-Mini-Instruct
No Fine-tuning GenRead 0.376 0.331 0.567 0.442 0.840 0.511
Always RAG, Wiki 0.460 0.541 0.594 0.549 0.771 0.583
Always RAG, PubMed 0.676 0.143 0.356 0.447 0.726 0.470
SR-RAG (3-source) Always RAG, Wiki 0.486 0.580 0.634 0.676 0.805 0.636
Always RAG, PubMed 0.738 0.302 0.494 0.644 0.813 0.598
Always Verbalizing 0.498 0.327 0.510 0.704 0.819 0.572
SR-RAG 0.736 0.560 0.639 0.716 0.824 0.695
(% <Wiki>) 1% 95% 87% 54% 0% 47%
(% <Pubmed>) 97% 3% 10% 18% 27% 31%
(% <Self>) 2% 2% 3% 28% 73% 22%
Qwen2.5-7B-Instruct
No Fine-tuning GenRead 0.418 0.334 0.626 0.676 0.875 0.586
Always RAG, Wiki 0.612 0.563 0.667 0.446 0.916 0.641
Always RAG, PubMed 0.698 0.129 0.469 0.401 0.821 0.504
SR-RAG (3-source) Always RAG, Wiki 0.358 0.576 0.639 0.637 0.818 0.606
Always RAG, PubMed 0.726 0.307 0.503 0.676 0.824 0.607
Always Verbalizing 0.522 0.303 0.550 0.695 0.828 0.580
SR-RAG 0.730 0.572 0.651 0.681 0.834 0.694
(% <Wiki>) 15% 96% 86% 20% 0% 43%
(% <Pubmed>) 79% 1% 13% 48% 39% 36%
(% <Self>) 6% 3% 1% 32% 61% 21%
Table 6: Three-way routing between retrieving from Wikipedia, retrieving from PubMed, or self-verbalizing knowledge. For each model we report task accuracy and the selection frequencies of the individual sources wiki/PubMed/self under SR-RAG’s LLM+kNN knowledge source routing scheme.
C.2 Discussion: Further Applications
SR-RAG has numerous useful real-world applications and exhausitively experimenting on them is beyond the scope of this paper’s experiments. In this section, we discuss several potential applications that SR-RAG naturally supports.
Cost-Aware Routing
Querying different knowledge sources brings distinct latency cost. SR-RAG’s preference label can be constructed such that the cheapest useful knowledge source is preferred, improving the latency while preventing the performance degradation. In addition, varied levels of compute are required to verbalize the required knowledge for different types of queries. SR-RAG can be used to model this pattern by formulating multiple internal knowledge sources associated with different levels of compute. This philosophy connects with the literature on adaptive reasoning budget (Lou et al., 2025).
Recency-Aware Routing
Many user queries have an inherent temporal dimension (e.g., “today,” “as of Sept. 2025,” fast-moving entities) where stale knowledge degrades correctness. SR-RAG can expose “freshness-tiered” sources and learn a preference label that prioritizes the most up-to-date tier when temporal cues or out-of-vocabulary entities are detected. The router can condition on time expressions, named-entity recency features, and source timestamps. When the model’s self-estimate is uncertain or predates the query’s time scope, it escalates to a recency-guaranteed index.
Privacy-Aware Routing
Queries often mix public facts with sensitive tenant data (e.g., customer tickets, internal codenames, emails). SR-RAG can represent privacy tiers as distinct sources and train preference labels under a “least-privilege first” policy: prefer private indices when identifiers or confidentiality cues are present and fall back to public only when the query is non-sensitive and private stores are irrelevant.
Appendix D Further Analyses
D.1 Self-Routing RAG: Hyperparameters
In this section, we analyze other hyperparameters in SR-RAG.
Source Labeling Heuristics
In SR-RAG, the default design for source preference labeling is collecting the knowledge from each source and selecting the top-ranked source in top 50% knowledge contexts in terms of contribution to the likelihood of the answer. In this section, we compare this strategy to a number of alternative heuristics:
  • Best Single Likelihood: Selecting the source that produces the knowledge leading to the highest answer likelihood.
  • Best Average Likelihood: Selecting the source that leads to the highest answer likelihood, averaged over all knowledge contexts.
  • Best All Rank: Selecting top-ranked knowledge source using all the knowledge context instead of top-50%.
In Table˜7, we present the F1 score and AUROC using Llama-2-7B-Chat as the LLM. Correctness is defined as preferring self-verbalized knowledge when it is better than or equal to performing retrieval in terms of the downstream question answering performance. Compared to three baseline methods, the heuristics SR-RAG uses achieves the best F1 and AUROC over the short-form and closed-set subset of the training set. Building upon this paper’s results, future work can further study leveraging more advanced uncertainty quantification methods (Wu et al., 2024b; Moskvoretskii et al., 2025b) to build more accurate source selection labels.
Method ARC_Easy NQ OBQA Fever Average
F1 AUROC F1 AUROC F1 AUROC F1 AUROC F1 AUROC
Best Single Likelihood 0.640 0.572 0.320 0.551 0.660 0.606 0.560 0.650 0.545 0.585
Best Average Likelihood 0.730 0.692 0.420 0.626 0.720 0.643 0.630 0.703 0.630 0.666
Best All Rank 0.700 0.691 0.470 0.632 0.700 0.621 0.630 0.674 0.625 0.655
SR-RAG 0.740 0.691 0.440 0.635 0.720 0.635 0.670 0.711 0.643 0.668
Table 7: F1 and AUROC scores for different source labeling methods across datasets. Llama-2-7B-Chat is used as the LLM. The best score per column is boldfaced.
Figure 6: Hyperparameter settings for the kNN policy datastore with Llama-2-7B-Chat.
Datastore Size
To study the influence of the kNN policy datastore size on the performance of SR-RAG, we randomly sample a subset of the training set (50k samples in total) to construct the datastore. As shown in Figure˜6 (blue), reducing the kNN policy datastore to 25k (half) has minimal impact. Even with just 1k examples, performance drops only slightly, suggesting potential for lightweight and memory-efficient deployments.
Number of Neighbors
The number of neighbors retrieved from the datastore determines the localness of the focus in the hidden representation space. While task-specific neighborhood size can be tuned, SR-RAG generally works well with 10 to 30 neighboring data points (Figure˜6 middle), indicating its robustness. However, using a too small (e.g., 5) or a too large (e.g., more than 50) neighborhood harms the performance of SR-RAG.
Layer for Hidden State
We further study the impact of layer selection on SR-RAG’s performance in Figure˜6 part 3 (green). Overall, we observe that using a middle layer’s hidden states has more superior performance over using the last layer’s. In the next section, we visualize the hidden state space across LLMs and show that middle layers tend to learn better representations for the model uncertainty as well as task types.
D.2 Hidden State Space of the Source Reflection Token
Does SR-RAG training allow the model to condense their knowledge of task, model uncertainty, and source characteristics into the representation of <EOQ>?
Uncertainty Encoding in <EOQ>
In Figure˜9, we visualize <EOQ> hidden states on PopQA using t-SNE. Clear separation emerges between instances where self-verbalized knowledge is helpful (green) and unhelpful (red). Middle layers show stronger clustering than final layers, supporting the use of these representations for kNN-based source selection.
Implicit Task Clustering
We further randomly sample 500 data points from the SR-RAG train set and visualize their hidden states. Figure˜10 shows that SR-RAG also learns to cluster instances by task type (e.g., fact-checking, closed QA, long-form generation). This suggests that source reflection tokens carry semantically meaningful and task-sensitive representations on top of uncertainty information. These behaviors serve as the foundation of our nearest neighbor-based source selection approach.
Source for $c_{i+}$ and $c_{i-}$ PopQA TriviaQA PubHealth ARC Average
%Verb $\geq$ RAG $(\uparrow)$
Self-Generated Knowledge 33.0% 56.6% 73.4% 63.5% 56.6%
Distilled Knowledge 28.2% 55.3% 69.7% 62.3% 53.9%
Verbalization Performance $(\uparrow)$
Self-Generated Knowledge 0.302 0.572 0.734 0.634 0.561
Distilled Knowledge 0.282 0.553 0.697 0.606 0.535
Table 8: Comparison between training with self-generated knowledge versus knowledge distillation for the verbalization branch in SR-RAG. Llama-2-7B-Chat is fine-tuned either with self-verbalized knowledge or the knowledge verbalized from Llama-3.3-70B-Instruct.
D.3 Can SR-RAG learn more from self-generated knowledge?
In SR-RAG training, self-verbalized knowledge contexts from the LLM are utilized to fine-tune the model via behavior cloning (stage 1) or preference alignment (stage 2). Alternatively, is it possible that the model can learn to verbalize even higher-quality knowledge from off-policy samples generated by a stronger LLM? We investigate this hypothesis by running GenRead-based knowledge verbalization on Llama-3.3-70B-Instruct and directly use the knowledge to prepare SR-RAG training data with Llama-2-7B-Chat. Table˜8 compares the performance of fine-tuning Llama-2-7B-Chat with this data versus using self-generated knowledge for SR-RAG. Interestingly, using self-generated positive and negative knowledge pairs is more beneficial for unlocking the model’s ability to generate useful knowledge, outperforming the alternative by 4.9% higher answer accuracy and 5.0% higher win rate over retrieval. It is also possible that the ability of the teacher model needs to be close to the fine-tuned model for its knowledge to be useful. We leave this investigation to future work.
D.4 Qualitative Study
In Figure˜7 and Figure˜8, we show two qualitative examples from TriviaQA with the fine-tuned Llama-2-7B-Chat as the backbone LLM. For both examples, we show the LLM’s prediction conditioned on both retrieved knowledge and the verbalized knowledge. While the verbalized knowledge is more concise than retrieved paragraphs, it better identifies the correct answers for both questions. This is because the core information (i.e., "first name", "island") only takes up a small portion of the query. As a result, dense retrieval is noise-prone, bringing in only distantly relevant information. By contrast, the LLM as a knowledge generator dynamically adjusts its focus, guiding the knowledge generation to include the core information. As a result, the LLM can correctly answer both questions with self-verbalized knowledge but provides incorrect answers with the low-quality retrieved knowledge. In the source selection process, the LLM directly predicts the source preference $p_{M}(\text{{\textless s\textgreater}}|q)$. In addition, questions from the train set sharing the same topic and difficulty level are retrieved to form the source preference distribution $p_{D}(\text{{\textless s\textgreater}}|q)$. In the first example (Figure˜7), both distributions prefer the LLM itself as the knowledge source. By contrast, in the second example (Figure˜8), $p_{M}(\text{{\textless s\textgreater}}|q)$ has a high uncertainty and $p_{D}(\text{{\textless s\textgreater}}|q)$ corrects the source distribution to select the LLM itself as the knowledge source.
Question: In the book David Copperfield what was Mr. Micawber’s first name?
Answer: Wilkins
Retrieved Knowledge: [1] David Copperfield (1999 film) a visit from Mr. Peggotty, back from Australia. He brings news that Emily has made a full recovery and that Mr. Micawber has established himself as a successful magistrate and bank manager. The story closes with the birth of David and Agnes’s third child - a girl. Betsey Trotwood’s wish finally comes true after nearly 30 years, as David decides that the baby will be christened Betsey Trotwood Copperfield, in honour of her godmother. [2]… (following knowledge chunks omitted)
Prediction (with retrieved knowledge): John Micawber, Sr., is a fictional character in Charles Dickens’s 1850 novel David Copperfield. He is the father of David Copperfield’s godfather, Mr. Micawber, Jr., and is portrayed as a kind and gentle man who is often in financial difficulties. Mr. Micawber’s first name is Charles, not John, as is often mistaken.
Verbalized Knowledge: David Copperfield is a novel by Charles Dickens, published in 1849-50. The novel follows the life of David Copperfield, a young man who grows up in Victorian England. Mr. Micawber is a character in the novel, a kind and well-meaning man who is a friend of David’s father. Mr. Micawber’s first name is Wilkins.
Prediction (with verbalized knowledge): Wilkins Micawber. Mr. Micawber is a character in Charles Dickens ’s novel David Copperfield. He is a kind and well - meaning friend of David ’s father, Mr. Copperfield. Mr. Micawber is known for his optimistic outlook on life and his tendency to get into financial difficulties.
Neighbor Questions:
- What was Mr. Smith’s first name in Mr. Smith goes to Washington? (<Self> >> <Wiki>)
- Who was the original singer of Knocking on Heaven’s Door? (<Self> >> <Wiki>)
- When did sir Arthur Conan Doyle wrote his first book? (<Wiki> >> <Self>)
… (27 questions omitted)
$p_{D}(\text{{\textless s\textgreater}}|q)$: P(<Wiki>) = 9/30 = 0.3, P(<Self>) = 21/30 = 0.7.
$p_{M}(\text{{\textless s\textgreater}}|q)$: P(<Wiki>) = 0.146, P(<Self>) = 0.854.
Final Source Decision: <Self>
Figure 7: Example of SR-RAG inference where $p_{D}(\text{{\textless s\textgreater}}|q)$ agrees with $p_{M}(\text{{\textless s\textgreater}}|q)$. As the model itself is selected as the knowledge source, the retrieval will be abstained. The retrieved knowledge and the corresponding prediction are shown only for illustration.
Question: From 1791 to 1803, Toussaint L’Ouverture led a revolt against France on which island?
Answer: Haiti
Retrieved Knowledge: [1] Saint-Domingue expedition slave revolt that led to the abolition of slavery in 1793 by the civil commissioners Sonthonax and Polverel, in a decision endorsed and spread to all the French colonies by the National Convention 6 months later. Toussaint Louverture, a black former slave who had been made Governor by France, re-established peace, fought off Spanish and British attempts to capture the island, and reestablished prosperity by daring measures. [2]… (following knowledge chunks omitted)
Prediction (with retrieved knowledge): Saint-Pierre et Miquelon (Saint-Pierre and Miquelon), a French overseas collectivity in the North American Atlantic Ocean.
Verbalized Knowledge: Toussaint L’Ouverture was a Haitian leader who led a successful revolt against French colonial rule in Haiti from 1791 to 1803. L’Ouverture was a former slave who became a prominent figure in the Haitian Revolution, which was a series of uprisings and battles fought between Haitian slaves and their French colonial oppressors. After years of fighting, L’Ouverture and his forces were able to secure Haiti’s independence from France on January 1, 1804.
Prediction (with verbalized knowledge): Haiti (Saint-Domingue).
Neighbor Questions:
- The battle of Hastings in 1066 was fought in which country? (<Self> >> <Wiki>)
- When Belgium declared its independence in 1830 it broke away from ___ control? (<Self> >> <Wiki>)
- When did the French come to the new world? (<Self> >> <Wiki>)
… (27 questions omitted)
$p_{D}(\text{{\textless s\textgreater}}|q)$: P(<Wiki>) = 24/30 = 0.8, P(<Self>) = 6/30 = 0.2.
$p_{M}(\text{{\textless s\textgreater}}|q)$: P(<Wiki>) = 0.576, P(<Self>) = 0.424.
Final Source Decision: <Self>
Figure 8: Example of SR-RAG inference where $p_{D}(\text{{\textless s\textgreater}}|q)$ corrects the source selection from $p_{M}(\text{{\textless s\textgreater}}|q)$. The model itself is selected as the knowledge source.
Figure 9: Hidden states of <EOQ> across three LLMs on PopQA, visualized using t-SNE. Llama = Llama-2-7B-Chat, Phi = Phi-3.5-Mini-Instruct, and Qwen = Qwen2.5-7B-Instruct.
Figure 10: Hidden states of <EOQ> across three LLMs on a random sample of 500 data points in the SR-RAG training set, visualized using t-SNE. Llama = Llama-2-7B-Chat, Phi = Phi-3.5-Mini-Instruct, and Qwen = Qwen2.5-7B-Instruct.